### v1

# DG-DETR: Toward Domain Generalized Detection Transformer

Abstract
End-to-end Transformer-based detectors (DETRs) have demonstrated strong detection performance. However, domain generalization (DG) research has primarily focused on convolutional neural network (CNN)-based detectors, while paying little attention to enhancing the robustness of DETRs. In this letter, we introduce a Domain Generalized DEtection TRansformer (DG-DETR), a simple, effective, and plug-and-play method that improves out-of-distribution (OOD) robustness for DETRs. Specifically, we propose a novel domain-agnostic query selection strategy that removes domain-induced biases from object queries via orthogonal projection onto the instance-specific style space. Additionally, we leverage a wavelet decomposition to disentangle features into domain-invariant and domain-specific components, enabling synthesis of diverse latent styles while preserving the semantic features of objects. Experimental results validate the effectiveness of DG-DETR. Our code is available at https://github.com/smin-hwang/DG-DETR.
keywords:
Object detection, Domain generalization, Detection transformer, Query selection, Wavelet decomposition.
journal: Elsevier.
[…]
[a]organization=Artificial Intelligence Graduate School, Gwangju Institute of Science and Technology (GIST), city=Gwangju, postcode=61005, country=Republic of Korea
[…]
[b]organization=School of Electrical Engineering and Computer Science, Gwangju Institute of Science and Technology (GIST), city=Gwangju, postcode=61005, country=Republic of Korea
1 Introduction
Deep neural network (DNN)-based object detectors [1, 2, 3, 4] have shown remarkable performance in various computer vision tasks under the independent and identically distributed (i.i.d.) assumption that training and test data are sampled from the same distribution. However, when these models are deployed in real-world scenarios where the distribution differs from the training set, their performance often degrades significantly due to domain-shifts [5, 6, 7]. This issue is especially critical in safety-sensitive applications such as autonomous driving. In this regard, addressing the distribution shifts between source and target domains has remained a long-standing challenge in the computer vision community.
To alleviate this problem, the prominent research direction is unsupervised domain adaptation (UDA) [8, 9, 10, 11, 12, 13], which aims to align the data distribution of a labeled source domain to that of an unlabeled target domain. Although UDA methods have demonstrated promising results, they heavily rely on the assumption that target domain data is available during the adaptation process. Moreover, collecting sufficient training data for all target domains of interest is often impractical [14] due to the cost and effort required, even without annotations.
Domain generalization (DG) [15, 16, 17, 14] has emerged as a feasible solution, aiming to train a model that generalizes well to unseen target domains by learning from multiple observed source domains during training. Most DG methods generally strive to learn domain-invariant feature representations across these source domains, but their performance is sensitive to the diversity and quantity of the available source domains [18, 19], which are often costly and labor-intensive to collect. As a more practical approach, recent research has begun focusing on single-domain generalization (S-DG) [20, 21], which aims to train robust models using only one source domain. However, achieving OOD robustness in this setting remains an open challenge.
In recent years, research on S-DG for object detection [22, 23, 24] has been limited, despite its critical importance in many visual perceptual systems. Furthermore, existing S-DG studies have largely focused on CNN-based detectors [1, 2, 3], although Vision Transformers (ViTs) have demonstrated strong generalization capabilities attributed to their superior shape bias, which stems from their ability to capture global dependencies among tokens [25, 26]. Building on this, in this letter, we explore the potential of Detection Transformer (DETR) [4], which shares the global modeling advantage, for single-domain generalization in object detection.
Figure 1: Overview of the proposed DG-DETR framework.
Similar to the human visual system, which exhibits high robustness by extracting global shape information [27], recent studies [28, 29, 30] have shown that shape-biased models achieve high robustness. Inspired by this observation, we aim to increase shape bias both implicitly and explicitly to improve model robustness against domain shifts. In this work, we propose a domain-agnostic query selection strategy to explicitly remove domain-induced biases from queries and select those that contain rich domain-invariant representations (e.g., shape). This approach provides high-quality initial queries to the decoder for robust object detection.
To implicitly improve model robustness, style augmentation [31, 32], which assumes that domain shifts are caused by style variance, has been successfully applied to single-domain generalization in image classification, generating multiple domains from a single source domain. However, style synthesis may potentially distort image contents, which are essential for the hierarchical object detection where a diverse context is present. Therefore, the key challenge in applying style augmentation to object detection lies in preserving the shape and semantics of objects while achieving diverse style augmentation. To do this, we utilize wavelet transform to decompose the domain-invariant features and domain-specific features. This approach enables us to perturb only the domain-specific features (e.g., texture), while preserving the content structures (e.g., shape) inherent in the images. The key contributions of this letter are summarized as follows:
  • 1.
    To the best of our knowledge, this is the first study to explore the potential of DETRs for domain generalized object detection.
  • 2.
    We introduce DG-DETR, a simple yet effective plug-and-play method that improves the generalization capability of DETRs.
  • 3.
    We propose a domain-agnostic query selection strategy that explicitly removes domain-induced biases from object queries.
  • 4.
    To preserve the semantics of objects during style augmentation, we leverage wavelet decomposition to perturb only domain-specific features.
2 Proposed Method
In this section, we briefly introduce the overall architecture of DG-DETR. We then describe the newly introduced components for robust object detection in detail.
2.1 Overview
DETRs consist of a CNN backbone, an encoder-decoder structure, and object class and box position predictors. To improve generalization capability, we introduce a style augmentation module and domain-agnostic query selection. The overall framework of DG-DETR is illustrated in Fig. 1.
Given an input image, the backbone network first extracts image features, while the style augmentation module synthesizes diverse latent styles. The output is then fed into an efficient hybrid encoder [33] which combines Transformer and CNN to enhance feature representation. The encoded features are subsequently passed to the domain-agnostic query selection module. This module removes domain-induced biases from queries and provides high-quality initial queries for the decoder. Finally, the decoder with auxiliary prediction heads iteratively refines object queries to generate object categories and bounding boxes. We note that while RT-DETR [33] is adopted as an example, the DG-DETR framework is compatible with other DETR-based detectors.
Figure 2: Visualization of wavelet decomposition. The image is decomposed into low- and high-frequency components, where low-frequency bands capture overall style and high-frequency bands preserve object shapes.
Figure 3: Overview of our style augmentation module. Wavelet transform decomposes the shallow CNN features into low- and high-frequency components. We perturb only the low-frequency features components.
2.2 Wavelet-guided Style Augmentation Module
Augmenting styles potentially cause the problem of semantic drift in objects, leading to imprecise object localization and misclassification. Inspired by the properties of the frequency spectrum [34, 26], we argue that frequency-based style augmentation enables the synthesis of diverse domains while preserving object semantics. Recall that the high-frequency components capture more global features (e.g., shape) while the low-frequency components contain smooth surfaces and textures (see Fig. 2). This suggests that applying style augmentation only to the low-frequency components affects overall texture while preserving the image contents.
To this end, we propose to utilize wavelet transform to disentangle high- and low-frequency features of images. Specifically, given an intermediate CNN feature map $F\in\mathbb{R}^{H\times W\times C}$ with spatial dimensions $H\times W$ and $C$ channels, discrete wavelet transform (DWT) applies four kernels, $LL^{T},LH^{T},HL^{T}$, and $HH^{T}$, where the low-pass filter is defined as $L=\left[\begin{array}[]{ll}1/\sqrt{2}&1/\sqrt{2}\end{array}\right]$ and high pass filter is $H=\left[\begin{array}[]{ll}-1/\sqrt{2}&1/\sqrt{2}\end{array}\right]$. These kernels slide across the input $F$ to decompose it into the four wavelet sub-bands: $F_{ll},F_{lh},F_{hl},$ and $F_{hh}$, each with reduced resolution. In this letter, we denote the low-frequency components $F_{ll}$ as $F_{low}$ and the high-frequency components $[F_{lh},F_{hl},F_{hh}]$ as $F_{high}$, respectively.
As shallow CNN layers preserve more style information [35], we apply the style augmentation module to the CNN backbone. Specifically, as a common practice, we adopt the feature channel statistics (i.e., mean and variance) to represent image styles. We then perturb the style statistics of source domain training instances to synthesize novel domain styles. Several prior studies [32, 31, 36, 23] have explored domain synthesis in feature space by perturbing feature statistics. To balance diversity and fidelity, we employ Normalization Perturbation (NP) [23] to perturb feature statistics in the shallow CNN backbone layers.
Given $F_{ll}$, we insert random noise into the feature channel statistics using Normalization Perturbation (NP) as follows:
$\hat{F_{ll}}=\mathrm{NP}(F_{ll}).$ (1)
The Normalization Perturbation (NP) is formulated as:
$y=\sigma_{s}^{*}\frac{x-\mu_{c}}{\sigma_{c}}+\mu_{s}^{*},$ (2)
$\sigma_{s}^{*}=\alpha\sigma_{c},$ (3)
$\mu_{s}^{*}=\beta\mu_{c},$ (4)
where $\{\mu_{c},\sigma_{c}\}\in\mathbb{R}^{C}\text{ and }\{\mu_{s},\sigma_{s}\}\in\mathbb{R}^{C}$ represent the mean and variance of the input content image and the stylized image, respectively.
Afterward, $\hat{F_{ll}}$ and $F_{high}$ are fed into the inverse discrete wavelet transform (IDWT) layer to reconstruct the style-augmented feature map $\hat{F}\in\mathbb{R}^{H\times W\times C}$. The overall process is illustrated in Fig. 3. In this letter, we refer to our style augmentation module as WaveNP. Following [23], WaveNP is applied at stages 1 and 2 of the CNN backbone during model training.
2.3 Domain-agnostic Query Selection
In DETR models, object queries are a fixed set of learnable embeddings that serve as input to the Transformer decoder. These queries serve as latent object representations that interact with global image features in the decoder and are progressively optimized to map to object instances in images. Since optimizing learnable object queries is inherently challenging, several methods [37, 38] have introduced query selection schemes that use confidence scores to select the top-K features as initial object queries. Domain shifts, which primarily comes from variations in visual style, introduce significant style-induced biases (i.e., domain-induced biases) into latent features, thereby limiting the generalization ability of the learned model. Therefore, removing domain-induced biases from queries enhances the robustness of DETR models against real-world domain shifts.
[…]
$s=E_{s}(\mu_{s}+\sigma_{s}),$ (5)
where $s\in\mathbb{R}^{D}$ is a style embedding, $D$ is the feature dimension, and $E_{s}$ is a style encoder consisting of a linear layer and a normalization layer.
Let $L$ represent the subspace spanned by $s$ and $Q\in\mathbb{R}^{N\times D}$, where $N$ is the number of feature sequences, denote the flattened encoded image features. We aim to remove the component of $Q$ that lies along the axis of $L$:
$\hat{Q}=Q-\alpha\text{Proj}_{L}Q,$ (6)
where $\alpha$ is a hyperparameter between 0 and 1 that controls the degree of removal of the style components. Note that during training, $\alpha$ is fixed at 1.
From $\hat{Q}$, we use confidence scores to select the top-K scoring features as the initial object queries:
$\hat{Q}_{select}=\text{Top-}K(E_{c}(\hat{Q})),$ (7)
where $\hat{Q}_{select}$ represents the set of K selected features, and $E_{c}$ is an auxiliary prediction head to select the top-K features.
We note that removing style-induced biases from image features may incur the loss of meaningful information. Therefore we subtract style-relevant components from feature sequences only in query selection process.
Figure 4: Visualization of the orthogonal projection. The query $Q$ is projected onto style-irrelevant semantic axes.
Table 1: Quantitative comparison of domain generalization performance on the Diverse Weather Dataset (DWD). The results represent mAP (%) for different weather conditions when trained on Daytime-Sunny and tested on Night-Sunny, Dusk-Rainy, Night-Rainy, and Daytime-Foggy.
Night-Sunny Dusk-Rainy Method bus bike car motor person rider truck mAP Method bus bike car motor person rider truck mAP RT-DETR [33] 45.4 41.2 70.5 19.1 48.9 30.8 47.6 43.3 RT-DETR [33] 46.1 23.7 74.0 16.1 36.7 16.3 49.7 37.5 SW [39] 34.6 23.7 65.7 10.2 37.1 19.7 35.2 32.3 SW [39] 40.0 14.2 72.6 6.8 34.7 10.5 41.6 31.5 IBN-Net [35] 36.8 24.0 65.5 13.1 39.9 21.2 38.9 34.2 IBN-Net [35] 43.2 16.0 72.3 14.4 35.4 12.2 44.9 34.1 IterNorm [40] 36.8 31.9 68.4 14.6 44.9 25.7 38.8 37.3 IterNorm [40] 35.4 17.4 69.1 15.7 29.7 11.9 39.7 31.3 NP [23] 49.0 40.5 71.9 20.8 49.2 33.4 50.2 45.0 NP [23] 51.8 28.6 77.4 19.0 42.2 22.3 54.1 42.2 DG-DETR (ours) 51.2 44.7 73.4 23.0 51.9 36.1 52.5 47.6 DG-DETR (ours) 51.0 28.1 77.7 20.1 42.0 19.9 55.7 42.1
Night-Rainy Daytime-Foggy Method bus bike car motor person rider truck mAP Method bus bike car motor person rider truck mAP RT-DETR [33] 27.7 7.8 42.0 0.3 12.9 8.7 24.8 17.7 RT-DETR [33] 32.7 27.5 61.8 30.6 38.1 36.9 25.0 36.1 SW [39] 20.3 6.5 42.9 0.1 10.8 4.1 18.7 14.8 SW [39] 23.5 21.4 55.7 20.4 31.1 32.7 12.9 28.2 IBN-Net [35] 29.3 6.6 42.6 0.9 13.7 8.0 24.9 18.0 IBN-Net [35] 19.9 17.8 47.2 19.3 25.7 24.1 13.0 23.9 IterNorm [40] 18.6 7.4 33.6 0.6 9.1 7.9 12.8 12.9 IterNorm [40] 25.9 25.5 59.2 24.7 37.4 36.1 15.6 32.1 NP [23] 39.9 16.8 50.4 1.0 17.7 8.9 32.0 23.9 NP [23] 33.2 27.1 61.7 29.2 38.3 37.6 25.9 36.2 DG-DETR (ours) 39.1 19.6 51.3 4.5 18.3 14.0 32.7 25.6 DG-DETR (ours) 36.4 28.9 64.4 31.2 40.3 39.9 28.8 38.5
Table 2: In-domain performance comparison on the Daytime-Sunny of the Diverse Weather Dataset (DWD). All models are trained and tested under the same weather condition.
Daytime-Sunny Method bus bike car motor person rider truck mAP RT-DETR [33] 61.3 48.3 82.2 51.7 59.8 49.2 63.1 59.4 SW [39] 55.4 40.4 80.5 42.9 54.5 44.5 55.9 53.4 IBN-Net [35] 56.0 37.3 80.0 42.7 53.9 43.1 57.3 52.9 IterNorm [40] 60.4 47.3 82.4 46.2 59.7 47.9 61.2 57.9 NP [23] 61.1 49.7 82.9 49.6 60.5 51.9 63.4 59.9 DG-DETR (ours) 62.5 50.7 83.5 51.9 61.6 53.3 64.9 61.2
Table 3: Ablation analysis of the proposed components. The results presents the impact of Wavelet-guided Style Augmentation (WaveNP) and Domain-agnostic Query Selection (DAQS) on domain generalization performance.
WaveNP DAQS DS NS DR NR DF
59.4 43.3 37.5 17.7 36.1
60.7 46.7 42.4 25.0 37.7
60.6 46.3 39.2 20.9 38.4
61.2 47.6 42.1 25.6 38.5
Table 4: Ablation analysis on perturbation of different frequency components. We compares the impact of applying feature perturbation to high-frequency vs. low-frequency components on domain generalization performance.
High Low DS NS DR NR DF
59.3 43.8 37.8 18.5 36.2
60.7 46.7 42.4 25.0 37.7
3 Experiments
In this section, we evaluate the generalization capability of our method against out-of-distribution scenarios including an ablation study to verify the effectiveness of proposed components.
3.1 Experimental setup
To evaluate our method, we use the Diverse Weather Dataset (DWD)[22], an urban-scene detection benchmark that includes five different weather conditions: Daytime-Sunny (DS), Night-Sunny (NS), Night-Rainy (NR), Dusk-Rainy (DR), and Daytime-Foggy (DF). DWD collected images from BDD-100k [41], FoggyCityscapes [7], and AdverseWeather [42] datasets. Following [22], we train the model using only the source domain (i.e., Daytime-Sunny) and evaluate it directly on the other adverse weather domains.
In all experiments, we adopt RT-DETR [33] as the base detector. To the best of our knowledge, no prior study has explored domain generalization (DG) in DETR-based object detection. Therefore, we extend existing DG methods [39, 35, 40, 18, 23], originally implemented for CNNs, to RT-DETR. All these methods improve the generalization ability of models by applying feature normalization  [39, 35, 40, 18] or perturbation [23] to the CNN feature extractor. Thus, we directly integrate them into DETR’s feature extractor (i.e., CNN backbone). To evaluate domain generalization performance, we follow [22] and use the Mean Average Precision (mAP) metric. Specifically, we report mAP at an Intersection over Union (IoU) threshold of 0.5 (mAP@0.5). The projection scaling factor (Eq. 6) is set to 1.0.
Figure 5: Detection results on Night-Sunny, Dusk-Rainy, Night-Rainy, Daytime-Foggy. First Row: RT-DETR [33] results. Second Row: The results with our method.
3.2 Performance Analysis of Domain Generalization
Table 1 presents the domain generalization (DG) performance under real-world weather conditions, each representing different domain shift scenarios. Our method achieves the best performance on three of the datasets. Specifically, compared with the baseline, DG-DETR improves performance by 4.3%, 4.6%, 7.9%, and 2.4% on Night-Sunny (NS), Dusk-Rainy (DR), Night-Rainy (NR), and Daytime-Foggy (DF), respectively. On the other hand, normalization-based methods[39, 35, 40, 18] fail to improve out-of-distribution (OOD) robustness. This is because these methods negatively affect the feature discrimination ability, thereby weakening detection performance, as reported in[22].
We also conducted comparative experiments in in-domain scenarios, and the results are reported in Table 2. The results show that DG-DETR significantly outperforms existing methods, demonstrating that the proposed approach improves performance even when the training and test sets come from the same domain.
3.3 Ablation Analysis
To further evaluate the effectiveness of our key modules and strategies in DG-DETR, we conduct a series of ablation experiments. Table 3 presents the ablation study on each component. As observed, removing any single component degrades DG performance, demonstrating the contribution of each module. We note that Domain-agnostic Query Selection (DAQS) improves model performance even without augmentation or normalization techniques by removing domain-induced biases from object queries.
Table 4 presents an ablation analysis on the effect of feature perturbation to different frequency components. Perturbing only high-frequency components results in lower performance across all weather conditions. In contrast, perturbing low-frequency components leads to performance improvement, demonstrating that low-frequency perturbation effectively improves generalization ability by learning domain-invariant representations.
3.4 Qualitative Results
We present qualitative comparisons across different weather scenarios in Fig. 5. As observed, the baseline RT-DETR fails to detect objects in challenging environments, often missing detections or producing inaccurate predictions. In contrast, our method shows more precise and reliable object detection under adverse conditions, demonstrating its robustness in cross-domain scenarios.
4 Conclusion
In this letter, we propose Domain Generalized DEtection TRansformer (DG-DETR) for DETR-based single-domain generalization. WaveNP preserves object semantics during feature perturbation by disentangling domain-invariant and domain-specific features. Domain-agnostic query selection improves the OOD robustness by removing domain-induced biases from object queries. Experimental results demonstrate that DG-DETR significantly improves the generalization ability of DETR models to unseen domains.
Acknowledgments